OpenStudy (anonymous):

What is the value of y so that the ordered pair (5, y) is a solution to the equation –5x + y = –40? explain it to me please ^-^ A. –65 B. 9 C. –15 D. 7

OpenStudy (anonymous):

5x+y=-40; (5,y) Plug in (5) into the equation. 5(5)+y=-40 25+y=-40 y=-40-25 y=-65 Any questions? :)

OpenStudy (anonymous):

It should be A. -65 So if you want to check your answer you can do it like, (5, -65) Plug this point into your original equation. 5x+y=-40 5(5)+(-65)=-40 25-65=-40 -40=-40 therefore, (x,y) are correct solutions.

OpenStudy (mathstudent55):

@karinaGperez Try it this way: –5x + y = –40 Plug in 5 for x and solve for y -5(5) + y = -40 -25 + y = -40 y = -15